Find a number such that the square of 1 more than the number is four more than four times the number.

Answer by Alan3354(69443)

Find a number such that the square of 1 more than the number is four more than four times the number.
------------------
(n+1)^2 = 4n + 4
n^2 + 2n + 1 = 4n + 4
n^2 -2n - 3 = 0
(n-3)*(n+1) = 0
n = 3, -1